$SPAD/src/input odpol.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool odpol.output
)set message test on
)set message auto off
)clear all
--S 1 of 36
dpol:= ODPOL(FRAC INT)
--R
--R
--R   (1)  OrderlyDifferentialPolynomial(Fraction(Integer))
--R                                                                 Type: Domain
--E 1
--S 2 of 36
w := makeVariable('w)$dpol
--R
--R
--R   (2)  theMap(DPOLCAT-;makeVariable;AM;17!0,0)
--RType: (NonNegativeInteger -> OrderlyDifferentialPolynomial(Fraction(Integer)))
--E 2
--S 3 of 36
z := makeVariable('z)$dpol
--R
--R
--R   (3)  theMap(DPOLCAT-;makeVariable;AM;17!0,0)
--RType: (NonNegativeInteger -> OrderlyDifferentialPolynomial(Fraction(Integer)))
--E 3
--S 4 of 36
w.5
--R
--R
--R   (4)  w
--R         5
--R                       Type: OrderlyDifferentialPolynomial(Fraction(Integer))
--E 4
--S 5 of 36
w 0

--R
--R
--R   (5)  w
--R                       Type: OrderlyDifferentialPolynomial(Fraction(Integer))
--E 5
--S 6 of 36
[z.i for i in 1..5]
--R
--R
--R   (6)  [z ,z ,z ,z ,z ]
--R          1  2  3  4  5
--R                 Type: List(OrderlyDifferentialPolynomial(Fraction(Integer)))
--E 6
--S 7 of 36
f:= w.4 - w.1 * w.1 * z.3
--R
--R
--R               2
--R   (7)  w  - w  z
--R         4    1  3
--R                       Type: OrderlyDifferentialPolynomial(Fraction(Integer))
--E 7
--S 8 of 36
g:=(z.1)**3 * (z.2)**2 - w.2
--R
--R
--R          3  2
--R   (8)  z  z   - w
--R         1  2     2
--R                       Type: OrderlyDifferentialPolynomial(Fraction(Integer))
--E 8
--S 9 of 36
D(f)
--R
--R
--R               2
--R   (9)  w  - w  z  - 2w w z
--R         5    1  4     1 2 3
--R                       Type: OrderlyDifferentialPolynomial(Fraction(Integer))
--E 9
--S 10 of 36
D(f,4)
--R
--R
--R   (10)

--R            2                               2
--R     w  - w  z  - 8w w z  + (- 12w w  - 12w  )z  - 2w z w
--R      8    1  7     1 2 6         1 3      2   5     1 3 5
--R   +
--R                                         2
--R     (- 8w w  - 24w w )z  - 8w z w  - 6w  z
--R          1 4      2 3  4     2 3 4     3  3
--R                       Type: OrderlyDifferentialPolynomial(Fraction(Integer))
--E 10
--S 11 of 36
df:=makeVariable(f)$dpol
--R
--R
--R   (11)  theMap(DPOLCAT-;makeVariable;AM;17!0,0)
--RType: (NonNegativeInteger -> OrderlyDifferentialPolynomial(Fraction(Integer)))
--E 11
--S 12 of 36
df.4
--R
--R
--R   (12)
--R            2                               2
--R     w  - w  z  - 8w w z  + (- 12w w  - 12w  )z  - 2w z w
--R      8    1  7     1 2 6         1 3      2   5     1 3 5
--R   +
--R                                         2
--R     (- 8w w  - 24w w )z  - 8w z w  - 6w  z
--R          1 4      2 3  4     2 3 4     3  3
--R                       Type: OrderlyDifferentialPolynomial(Fraction(Integer))
--E 12
--S 13 of 36
order(g)
--R
--R
--R   (13)  2
--R                                                        Type: PositiveInteger
--E 13
--S 14 of 36
order(g, 'w)
--R
--R
--R   (14)  2
--R                                                        Type: PositiveInteger
--E 14
--S 15 of 36

differentialVariables(g)
--R
--R
--R   (15)  [z,w]
--R                                                           Type: List(Symbol)
--E 15
--S 16 of 36
degree(g)
--R
--R
--R           2  3
--R   (16)  z  z
--R          2  1
--R                  Type: IndexedExponents(OrderlyDifferentialVariable(Symbol))
--E 16
--S 17 of 36
degree(g, 'w)
--R
--R
--R   (17)  1
--R                                                        Type: PositiveInteger
--E 17
--S 18 of 36
weights(g)
--R
--R
--R   (18)  [7,2]
--R                                               Type: List(NonNegativeInteger)
--E 18
--S 19 of 36
weights(g,'w)
--R
--R
--R   (19)  [2]
--R                                               Type: List(NonNegativeInteger)
--E 19
--S 20 of 36
weight(g)
--R
--R
--R   (20)  7
--R                                                        Type: PositiveInteger
--E 20
--S 21 of 36

isobaric?(g)
--R
--R
--R   (21)  false
--R                                                                Type: Boolean
--E 21
--S 22 of 36
eval(g,['w::Symbol],[f])
--R
--R
--R                  2                           2        3  2
--R   (22)  - w  + w  z  + 4w w z  + (2w w  + 2w  )z  + z  z
--R            6    1  5     1 2 4      1 3     2   3    1  2
--R                       Type: OrderlyDifferentialPolynomial(Fraction(Integer))
--E 22
--S 23 of 36
eval(g,variables(w.0),[f])
--R
--R
--R           3  2
--R   (23)  z  z   - w
--R          1  2     2
--R                       Type: OrderlyDifferentialPolynomial(Fraction(Integer))
--E 23
--S 24 of 36
monomials(g)
--R
--R
--R            3  2
--R   (24)  [z  z  ,- w ]
--R           1  2     2
--R                 Type: List(OrderlyDifferentialPolynomial(Fraction(Integer)))
--E 24
--S 25 of 36
variables(g)
--R
--R
--R   (25)  [z ,w ,z ]
--R           2  2  1
--R                              Type: List(OrderlyDifferentialVariable(Symbol))
--E 25
--S 26 of 36
gcd(f,g)
--R
--R

--R   (26)  1
--R                       Type: OrderlyDifferentialPolynomial(Fraction(Integer))
--E 26
--S 27 of 36
groebner([f,g])
--R
--R
--R                 2     3  2
--R   (27)  [w  - w  z ,z  z   - w ]
--R           4    1  3  1  2     2
--R                 Type: List(OrderlyDifferentialPolynomial(Fraction(Integer)))
--E 27
--S 28 of 36
lg:=leader(g)
--R
--R
--R   (28)  z
--R          2
--R                                    Type: OrderlyDifferentialVariable(Symbol)
--E 28
--S 29 of 36
sg:=separant(g)
--R
--R
--R            3
--R   (29)  2z  z
--R           1  2
--R                       Type: OrderlyDifferentialPolynomial(Fraction(Integer))
--E 29
--S 30 of 36
ig:=initial(g)
--R
--R
--R           3
--R   (30)  z
--R          1
--R                       Type: OrderlyDifferentialPolynomial(Fraction(Integer))
--E 30
--S 31 of 36
g1 := D g
--R
--R
--R            3               2  3
--R   (31)  2z  z z  - w  + 3z  z
--R           1  2 3    3     1  2

--R                       Type: OrderlyDifferentialPolynomial(Fraction(Integer))
--E 31
--S 32 of 36
lg1:= leader g1
--R
--R
--R   (32)  z
--R          3
--R                                    Type: OrderlyDifferentialVariable(Symbol)
--E 32
--S 33 of 36
pdf:=D(f, lg1)
--R
--R
--R             2
--R   (33)  - w
--R            1
--R                       Type: OrderlyDifferentialPolynomial(Fraction(Integer))
--E 33
--S 34 of 36
prf:=sg * f- pdf * g1
--R
--R
--R            3         2        2  2  3
--R   (34)  2z  z w  - w  w  + 3w  z  z
--R           1  2 4    1  3     1  1  2
--R                       Type: OrderlyDifferentialPolynomial(Fraction(Integer))
--E 34
--S 35 of 36
lcf:=leadingCoefficient univariate(prf, lg)
--R
--R
--R            2  2
--R   (35)  3w  z
--R           1  1
--R                       Type: OrderlyDifferentialPolynomial(Fraction(Integer))
--E 35
--S 36 of 36
ig * prf - lcf * g * lg
--R
--R
--R            6         2  3        2  2
--R   (36)  2z  z w  - w  z  w  + 3w  z  w z
--R           1  2 4    1  1  3     1  1  2 2
--R                       Type: OrderlyDifferentialPolynomial(Fraction(Integer))

--E 36 )spool )lisp (bye) ———-

10

References^{[1] nothing}

11